Solve the system of equations using the substitution method. 3x-4y=17,
Answer by 10301650(32)

GIVEN: ---: 3x-4y=17
---: y=x-5
substitute Y into x-5
---: 3x-4(4x-5)=17
---: 3x-4x+20=17
---: -x=17-20
---: -x=-3
multiply both side with -1
---: -1(-x=-3)-1
---: x=3 (ANSWER)
.
.
.
then use the second given to find the answer for the Y
.
y=3-5
y=-2 (ANSWER)